OpenStudy (anonymous):

Algebra 2 quick question!?

OpenStudy (anonymous):

So i had to make this "pool" for a project. the dimensions are Length= 10, Width= 25, Depth= 8, my surface are is 810, and th evolume of water needed for pool is 1,875ft.^3 I have to make a "doggie pool " that is 1/4 these measurements. any ideas?

OpenStudy (anonymous):

you could just divide length,width, and depth all by 4 then recalculate all the dimensions

OpenStudy (anonymous):

taking 1/4 of all would make it 1/4^3 smaller

OpenStudy (anonymous):

you come to my rescue again! lol so... 10/4=2.5 25/4=6.25 8/4=2

OpenStudy (anonymous):

@completeidiot

OpenStudy (anonymous):

if you want volume, multiply those all together if you want lining, then look back at the previous question same for volume of water

OpenStudy (anonymous):

@completeidiot If it asks for 1/4 of all the measurements such as the volume and surface area...dividing all by 1/4 would make it 1/4^3=1/64 the volume of the bigger pool. if it's just asking about 1/4 the 2d size then you'd divide all by 4 but if you want everything to be 1/4 of the bigger ,you'd need to only divide one thing by 1/4 or you could divide two (Length, Width, and Depth) by 1/2 which would look semi decent as you could run the doggie pool along the length of the pool

OpenStudy (anonymous):

@Outkast3r09 agreed for the first part for the 2nd part practically speaking, the dimensions of a rectangular prism pool being 10x24x2 or 2.5x24x8 etc the size would be a bit ridiculous for a dog, as the dimensions are all in ft

OpenStudy (anonymous):

it really depends on the dog

OpenStudy (anonymous):

yes 2.5 ft is pretty crazy but if you divide two by 1/2 say the width and the depth youd get width =12.5 length =10 depth=4

OpenStudy (anonymous):

@completeidiot thankyou for all your help. just one more thing! i dont know how to answer- Was the amount of pool liner material representative of the lateral or surface area of a rectangular prism? Why or why not? Im pretty sure it was surface area right

OpenStudy (anonymous):

or you could do 1/2 of length and depth width=25 length =5 depth =4

OpenStudy (anonymous):

@Outkast3r09 thankyou for all your help as well. I think i figuired this one out :)

OpenStudy (anonymous):